OpenStudy (anonymous):

checkin homework.....m-5 1/2 = 2 3/5

OpenStudy (anonymous):

\[5\frac{ 1 }{ 2 }=\frac{ 5*2+1 }{ 2 }=\frac{ 11 }{ 2}\] can you do the rest

OpenStudy (anonymous):

he got that part.....2 and 3/5 = 13 over 5

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now \[m-\frac{ 11 }{ 2 }=\frac{ 13 }{ 5 }\]

OpenStudy (anonymous):

\[m=\frac{ 11 }{ 2 }+\frac{ 13 }{ 5 }\]

OpenStudy (anonymous):

he came up with 8 and 1/10 asn the answer

OpenStudy (anonymous):

\[m=\frac{ 11 }{ 2 }+\frac{ 13 }{ 5 }=\frac{ 55+26 }{ 10 }=\frac{ 81 }{ 10 }=8\frac{ 1 }{ 10 }\]

OpenStudy (anonymous):

yea...thank you